OpenStudy (anonymous):

(HELP!!) ABCD is an isosceles trapezoid with diagonals that intersect at point P. If AB || CD, AC = 7y – 30, BD = 4y + 60, and CD = 5y + 14, solve for y. A.22 B.30 C.46 D.57

OpenStudy (anonymous):



ganeshie8 (ganeshie8):

hint : diagonals are congruent in isosceles trap

ganeshie8 (ganeshie8):

AC = BD

OpenStudy (anonymous):

7Y-30=4Y+60 7(30)-30=4(30)+60 180=180 ?

OpenStudy (anonymous):

B.

ganeshie8 (ganeshie8):

wat do u get for Y ?

OpenStudy (anonymous):

30

ganeshie8 (ganeshie8):

Excellent !

OpenStudy (anonymous):

THANKS! :D

ganeshie8 (ganeshie8):

YW ;)